Given the Production Function Q = 72X + 15X2 - X3, where Q =Output and X=Input (5 marks) a) What is the Marginal Product (MP) when X = 8? b) What is the Average Product (AP) when X = 6? c) At what value of X will Q be at its maximum? d) At what value of X will Diminishing Returns set in?


1
Expert's answer
2021-05-25T17:15:11-0400

Given "72X + 15X^2 - X^3"


a) "MP=\frac {dQ} {dX} =72+30X-3X^2"


When X=8;


"72+30(8)-3(8)^2 =72+240-192=120"


"MP=120"


b) "AP=\frac {Q} {X} =72+15X-X^2"


When X=6;


"AP=72+15(6)-(6\u00d76)=126"


"AP=126"


c) Q is maximum when "\frac {dQ} {dX} =0"


"72+30X-3X^2=0"


"3X^2-30X-72=0"


"X^2-10X-24=0"


"X^2-12X+2X-24=0"


"X(X-12)+2(X-12)=0"


"X=12" or (X="-2" which is inadmissible)


d) Diminishing returns means, MP is decreasing i.e

"\frac {dMP} {dX} \lt 0"

"\frac {dMP} {dX} =; 0"


"30-9X=0"


"9X=30"


"X=3.33"


Thus diminishing returns sets is when "X\gt 3.33"